# The Monge optimal transport barycenter problem

### v1

@@ 2 A Monge formulation of the optimal transport barycenter problem
Neither $\rho(x|z)$ (for any $z$) nor $\mu(y)$ assign finite measure to small sets. Then the fact that $T(:,z)$, the least costly map that pushes forward $\rho(:|z)$ to $\mu$, is invertible is a central result in optimal transport theory [3].
This theorem provides the ground for various applications.
@@
A third application addresses the following question: assuming that there exists a “true” additional source $w$ of variability in $x$, such that that $x=\phi(w,z)$, how much does our $y=T(x,z)$ teach us about the true $w$? Here the notion of a true source is field dependent. For our purposes, we just assume that such true $w$ exists.
An identifiability issue arises. In the absence of additional information, the conditional distribution $\rho(x|z)$ does not suffice to determine $w$. For instance, if for some value of $z$, $\phi\left(w_{1},z\right)=\phi\left(w_{2},z\right)$, then there is no way that using $x$ and $z$ alone we could distinguish between $w_{1}$ and $w_{2}$. More generally, any two $z$-dependent random variables $W_{1}^{z}$ and $W_{2}^{z}$ such that the distributions of both $\phi_{1}(z,W_{1}^{z})$ and $\phi_{2}(z,W_{2}^{z})$ agree with $\rho(x|z)$ explain the data equally well. In particular, $y$ provides one such explanatory variable, with the additional property that it is necessarily a function of $w$ and $z$:
$y=T(x,z)=T(\phi(w,z),z)=Y(w,z).$
Moreover, $Y(:,z)$ is invertible for all values of $z$ for which $w$ is identifiable, i.e. such that $\phi\left(w_{1},z\right)=\phi\left(w_{2},z\right)\Rightarrow w_{1}=w_{2}$, since
@@ 3 Adversarial characterization of independence
Posing the Monge optimal transport barycenter problem (2) in data-driven scenarios, where the joint distribution $\pi(x,z)$ is only known through $n$ sample pairs $\{x_{i},z_{i}\}$, requires a sample-friendly formulation of the independence condition between the random variables $y$ and $z$. We will use a weak characterization based on test functions [32, 5]: two variables $y\in\mathcal{Y}$ and $z\in\mathcal{Z}$ with joint distribution $\pi(y,z)$ are independent if and only if any two bounded measurable functions $g(y)$ and $f(z)$ satisfy
@@
We can restrict the functions $f$ and $g$ to smaller spaces $\mathcal{F}$ and $\mathcal{G}$, provided that they contain suitable approximations to the delta function consistent with the smoothness of $\pi(y,z)$. To prove this, notice that if $\pi\left(y,z\right)>\rho\left(y\right)\gamma\left(z\right)$ in a neighborhood $U_{*}$ of the pair $(y_{*},z_{*})$, then $g(y)=\delta_{y_{*}}\left(y\right)$, $f(z)=\delta_{z_{*}}\left(z\right)$ satisfy
$\mathbb{E}_{\pi}[g(y)f(z)]>\mathbb{E}_{\rho}[g(y)]\ \mathbb{E}_{\gamma}[f(z)],$
where $\delta_{y_{*}}$ and $\delta_{z_{*}}$ are non-negative functions whose product is positive at $(y_{*},z_{*})$ and vanishes outside of $U_{*}$.
We can further subtract from $f(z)$ its mean, yielding the following equivalence statement: two variables $y$ and $z$ are independent if and only if, for all functions $g(y)\in\mathcal{G}$ and $f(z)\in\mathcal{F}$ with $\mathbb{E}_{\gamma}[f(z)]=0$,
$\mathbb{E}_{\pi}[g(y)f(z)]=0.$
This equivalence gives rise to the following adversarial formulation of the barycenter problem (2):
@@
[…]
these norms would represent the standard deviation of $f$ and $g$, since not only $\mathbb{E}_{\gamma}[f]=0$ by construction, but also $\mathbb{E}_{\rho}[g]=0$ holds at the optimal solution: a constant added to $g$ does not affect the value of $L$, and the norm of $g-a$ is smallest when $a=\bar{g}$. It follows that we could read the problem as the minimization of the transportation cost subject to the condition that the correlation between any two functions $f(z)$ and $g(y)$ vanishes [32]. Yet adopting this choice for a norm is neither required nor convenient for $g(y)$, since it depends on the unknown $\rho(y)$, which evolves through the optimization procedure. We will propose an alternative norm below.
@@
A first requirement is to eliminate redundancy: since the columns of the matrix $G$ typically consist of smooth functions that we only evaluate at a finite set of points, the dimension of the effective range of $G$ can be much smaller than $m_{y}$, particularly when the latter is chosen large so as to accommodate for a rich set of candidate test functions. Eliminating such redundancy makes the maximization problem over $b$ smaller and better posed. The second requirement also relates to the ease of optimization over $b$: enforcing the requirement that $\|g(y)\|=1$ would be much easier if $G$ –an operator from $R^{m_{y}}$ to $\mathcal{G}$– were orthogonal, as it would directly translate into the condition that $\|b\|=1$.
Enforcing the orthogonality of $G$ requires that we fix an inner product in $\mathcal{G}$. The canonical one
$\phi,\psi\right>=\sum_{i}\phi\left(y_{i}\right)\psi\left(y_{i}\right),$ (7)
the empirical version of (5), has the problem that its very definition depends on the unknown $\{y_{i}\}$. We can stick to the canonical inner product in $z$-space, since the $\{z_{i}\}$ are fixed, but we should use a different one for functions of $y$. We need a functional norm such that a function $g(y)$ of norm $1$ cannot be very large on the data. As a counterexample, consider an inner product of the form
[…]
where the weighting function $w(y)$, though everywhere positive, is very small in at least one area where the true distribution $\rho(y)$ is not. Since the requirement that $\|g\|=1$ will effectively only constrain $g(y)$ in areas where $w(y)$ is comparatively large, the algorithm’s variables $b$ can make $L$ large by choosing test functions $g(y)$ not so much based on their correlation with $f(z)$ but just on their effective amplitude in areas where $\rho$ is large but the norm of $g$ does not truly act as a constraint. It follows that we must choose a weight $w(y)$ so that the Radon–Nikodym derivative $\frac{d\rho}{dw}$ is bounded. We adopt an inner product that is fixed through stages of the procedure, updating it only occasionally to reflect the evolving distribution of the $\{y_{i}\}$, replacing (7) by
$\phi,\psi\right>=\sum_{i}\phi\left(y^{0}_{i}\right)\psi\left(y^{0}_{i}\right),$ (8)
where $y^{0}_{i}$ is initially set to $x_{i}$ and then updated every so often, to reflect more accurately the different stages of the $\{y_{i}\}\sim\rho(y)$.
In order to replace $G(y)$ by a smaller dimensional, orthogonal operator $Q_{y}(y)$ that spans the same effective range, we perform the reduced singular-value decomposition
$G_{i}^{j}\stackrel{{\scriptstyle\text{def}}}{{=}}G^{j}\left(y^{0}_{i}\right)\approx\sum_{k=1}^{n_{y}}\sigma_{k}\ u^{k}_{i}\ v^{k}_{j},$
where $n_{y}\leq m_{y}$ is determined from the criterion that the sum $\sum_{k=1}^{n_{y}}\left(\sigma_{k}\right)^{2}$ be larger that a fraction of the squared norm of $\left\|G\right\|^{2}=\sum_{i,j}\left(G_{i}^{j}\right)^{2}=\sum_{k=1}^{m_{y}}\left(\sigma_{k}\right)^{2}$, and we adopt
$Q_{y}(y)=G(y)B^{y},\quad B^{y}_{jk}=\frac{1}{\sigma_{k}}v^{k}_{j}$
@@
(Notice that, in particular, $Q_{y}^{k}\left(y^{0}_{i}\right)=u^{k}_{i}$.) With a fixed inner product, the matrix $B^{y}$ needs to be computed only once per stage, when the $y^{0}$ are updated. The sense in which this $Q^{y}$ is orthogonal is not the conventional one: its columns represent the evaluation at arbitrary positions $\{y_{i}\}$ of a set of functions that are orthonormal under a chosen, fixed inner product based on the $\{y^{0}_{i}\}$.
@@ 4 A flow-based methodology
The maximization over $a$ and $b$ can be carried out explicitly: they must align with the left and right first principal components of the $n_{z}\times n_{y}$ matrix
$A^{hl}\left(y\right)\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i}Q^{h}_{z}\left(z_{i}\right)Q^{l}_{y}\left(y_{i}\right)$
($A$ depends only on the $\{y_{i}\}$, since the $\{z_{i}\}$ are fixed), and the penalty term is given by the first singular value $\sigma_{1}(y)$ of $A$. It follows that we can write the problem as a minimization over $y$ alone:
@@
$y_{i}^{n+1}=y_{i}^{n}-\eta^{n}\left[\frac{1}{n}\nabla_{y}\left.c\left(x_{i},y\right)\right|_{y_{i}^{n}}+\lambda\ a^{\prime}\left.\nabla_{y}A\right|_{y_{i}^{n}}b\right]$ (10)
(for which all $\{y_{i}\}$ decouple), where
$\left.\nabla_{y}A^{hl}\right|_{y_{i}^{n}}=Q^{h}_{z}\left(z_{i}\right)\sum_{j}\left.\nabla G^{j}(y)\right|_{y_{i}}B^{y}_{jl}$
and $a$ and $b$ are updated in an alternate step.
@@
where $a$ and $b$ are the left and right $k$th principal components of $A$.
@@
The penalty term $\sigma_{1}(y)$ is not smooth at its arg-min $y=y^{*}$: for $y$ and $z$ to be independent, the first singular value of $A$ must vanish, so $A(y^{*})$ itself must equal zero, and the first singular value $\sigma_{1}(y)$ of a matrix that depends smoothly on $y$ typically has corners where $A(y)$ vanishes (The simplest example is the $1\times 1$ matrix $A=y\in R$, whose only singular value $\sigma=|y|$ has a corner at $y=0$.) To address this, we square the penalty term:
$\min_{\{y_{i}\}}L=\sum_{i=1}^{n}c\left(x_{i},y_{i}\right)+\lambda\ \sigma_{1}^{2}(y),\quad\sigma_{1}(y)\stackrel{{\scriptstyle\text{def}}}{{=}}\max_{\|a\|=\|b\|=1}a^{\prime}A(y)b.$
There still remains one issue to address to make the methodology fully functional. Because every step of the algorithm brings down the largest singular value of $A(y)$, the first few of those singular values will tend to coalesce at convergence at a common value $\sigma_{*}\ll 1$. Then the derivatives of the penalty term with respect to the $\{y_{i}\}$ are not well defined, as they depend on the arbitrary choice of one pair among the various singular components $(a,b)$ associated to the singular value $\sigma_{*}$. In terms of test functions, more that one pair of functions $(f,g)$ have reached the threshold correlation $\sigma_{*}$.
To address this, we modify the algorithm so that it tracks the first $K$ pairs $(a_{k},b_{k})$ of principal component of $A$, where $K=\min(\hbox{rank}(A),K_{max})$, with $K_{max}$ fixed by the user. Then we descend over $y$
@@
Notice that this extension carries little computational cost, since the $A(y)$ to differentiate is common to all the $\{\sigma_{k}\}$. If performed using reproducing Kernel Hilbert spaces, this extension could be thought as interpolating between COCO [26] and HSIC [33].
@@ 5 Map inversion
Since we do not know $T(x,z)$ in closed form, it could appear that we can only invert it by learning $X(y,z)$ from its $n$ available samples $\{x_{i},y_{i},z_{i}\}$, for instance through kernel regression, nearest neighbor or neural networks. Yet we can do much better than that and obtain a closed form for $T^{-1}$, exploiting the fact that the penalization parameter $\lambda$ is large but finite. Since at convergence the gradient $\nabla_{y}L$ is zero, or at least sufficiently small to satisfy a termination criterion, we have
$\nabla_{y_{i}}c\left(x_{i},y_{i}\right)+2\lambda\ \sum_{k=1}^{K}\sigma_{k}\nabla_{y}\sigma_{k}\Big{|}_{y_{i},z_{i}}=0,$
which for the canonical quadratic cost in (1) yields $x_{i}=y_{i}+2\lambda\ \sum_{k=1}^{K}\sigma_{k}\nabla_{y}\sigma_{k}\Big{|}_{y_{i},z_{i}}$. In order to invert the map for arbitrary values of $y$ and $z$, we extend the validity of this expression and write
@@
Finding $x^{*}_{i}=X(y_{i},z_{*})$ requires the values of $f_{k}(z_{*})$ and $\nabla g_{k}\big{|}_{y_{i}}$. Out of these, all the
$\left.\nabla g_{k}\right|_{y_{i}}=\left.\nabla_{y}G(y)\right|_{y_{i}}B^{y}b_{k}$
are already known, since they have been used at the last descent step. Hence we only need the $K$ numbers $\{f_{k}(z_{*})\}$, which we calculate by introducing a new row $F(z_{*})$ of $F$ and setting
$f_{k}(z_{*})=F(z_{*})B^{z}a_{k}.$
The inversion formula in (12) provides us with a valuable bonus: it shows that the dependence of $x$ on $z$ that our algorithm has uncovered is mediated by the $K$ functions $\{f_{k}(z)\}$, so we have inadvertently performed factor extraction. The extracted factors bring in insights about the mechanisms of the dependence of $x$ on $z$, while the gradient of these functions inform us of the sensitivity of $x$ with respect to changes in $z$.
When we consider density estimation in Section 8, it will be useful to notice that we have access not only to $X(y,z)$ but also to its derivatives,

### v2

@@ 2 A Monge formulation of the optimal transport barycenter problem
Since neither $\rho(x|z)$ (for any $z$) nor $\mu(y)$ assign finite measure to small sets, the fact that the least costly map $T(:,z)$ pushing forward $\rho(:|z)$ to $\mu$ is invertible is a central result in optimal transport theory [5].
This lemma provides the ground for various applications.
@@
A third application addresses the following question: assuming that there exists a “true” additional source $w$ of variability in $x$, such that that $x=\phi(w,z)$, how much does our $y=T(x,z)$ teach us about the true $w$? (Here the notion of a true source is field dependent; for our purposes, we just assume that such true $w$ exists.) An identifiability issue arises: without additional information, the conditional distribution $\rho(x|z)$ does not suffice to determine $w$. For instance, if for some value of $z$, $\phi\left(w_{1},z\right)=\phi\left(w_{2},z\right)$, then there is no way that using $x$ and $z$ alone we could distinguish between $w_{1}$ and $w_{2}$. More generally, any two $z$-dependent random variables $W_{1}^{z}$ and $W_{2}^{z}$ such that the distributions of both $\phi_{1}(W_{1}^{z},z)$ and $\phi_{2}(W_{2}^{z},z)$ agree with $\rho(x|z)$ explain the data equally well. In particular, $y$ provides one such explanatory variable, with the additional property that it is necessarily a function of $w$ and $z$: $y=T(x,z)=T(\phi(w,z),z)=Y(w,z)$. Moreover, $Y(:,z)$ is invertible for all values of $z$ for which $w$ is identifiable, i.e. such that $\phi\left(w_{1},z\right)=\phi\left(w_{2},z\right)\Rightarrow w_{1}=w_{2}$, since
@@ 3 Adversarial characterization of independence
Posing the Monge OTBP (1) in data-driven scenarios, where the joint distribution $\pi(x,z)$ is only known through $n$ sample pairs $\{x_{i},z_{i}\}$, requires a sample-friendly formulation of the independence condition between the random variables $y$ and $z$. We will use a weak characterization based on test functions [33, 7]: two variables $y\in\mathcal{Y}$ and $z\in\mathcal{Z}$ with joint distribution $\pi(y,z)$ are independent if and only if any two bounded measurable functions $g(y)$ and $f(z)$ satisfy
@@
We can restrict the functions $f$ and $g$ to smaller spaces $\mathcal{F}$ and $\mathcal{G}$, provided that they contain suitable approximations to the delta function. To prove this, notice that if $\pi\left(y,z\right)>\rho\left(y\right)\gamma\left(z\right)$ in a neighborhood $U_{*}$ of the pair $(y_{*},z_{*})$, then $g(y)=\delta_{y_{*}}\left(y\right)$, $f(z)=\delta_{z_{*}}\left(z\right)$ satisfy $\mathbb{E}_{\pi}[g(y)f(z)]>\mathbb{E}_{\rho}[g(y)]\ \mathbb{E}_{\gamma}[f(z)]$, where $\delta_{y_{*}}$ and $\delta_{z_{*}}$ are non-negative functions whose product is positive at $(y_{*},z_{*})$ and vanishes outside of $U_{*}$.
Substractig the mean of $f(z)$ yields the following equivalence statement: two variables $y$ and $z$ are independent if and only if, for all functions $g(y)\in\mathcal{G}$ and $f(z)\in\mathcal{F}$ with $\mathbb{E}_{\gamma}[f(z)]=0$, $\mathbb{E}_{\pi}[g(y)f(z)]=0$. This equivalence gives rise to the following adversarial formulation of the barycenter problem (1):
@@
$\left(f_{1},f_{2}\right)=\mathbb{E}\left[f_{1},f_{2}\right],\quad\left(g_{1},g_{2}\right)=\mathbb{E}\left[g_{1},g_{2}\right],$ (5)
these norms would represent the standard deviation of $f$ and $g$, since not only $\mathbb{E}_{\gamma}[f]=0$ by construction, but also $\mathbb{E}_{\rho}[g]=0$ holds at the optimal solution: a constant added to $g$ does not affect the value of $L$, and the norm of gag-a is smallest when $a=\bar{g}$. It follows that we could read the problem as the minimization of the transportation cost subject to the condition that the correlation between any two functions $f(z)$ and $g(y)$ vanishes [33]. Yet adopting this choice for a norm is neither required nor convenient for $g(y)$, since it depends on the unknown distribution $\rho$ underlying $y$, which evolves through the optimization procedure. We propose an alternative norm below.
@@
A first requirement is to eliminate redundancy: since the columns of the matrix $G$ typically consist of smooth functions that we only evaluate at a finite set of points, the dimension of the effective range of $G$ can be much smaller than $m_{y}$, particularly when the latter is chosen large so as to accommodate for a rich set of candidate test functions. Eliminating such redundancy makes the maximization problem over $b$ smaller and better posed. The second requirement also relates to the ease of optimization over $b$: enforcing the requirement that $\|g(y)\|=1$ would be much easier if $G$ –an operator from $R^{m_{y}}$ to $\mathcal{G}$– were orthogonal, as it would directly translate into the condition that $\|b\|=1$. Enforcing the orthogonality of $G$ requires that we fix an inner product in $\mathcal{G}$. The canonical one, $\left<\phi,\psi\right>=\sum_{i}\phi\left(y_{i}\right)\psi\left(y_{i}\right)$, the empirical version of (5), has the problem that its very definition depends on the unknown $\{y_{i}\}$. We can stick to the canonical inner product in $z$-space, since the $\{z_{i}\}$ are fixed, but we should use a different one for functions of $y$. We need a functional norm such that a function $g(y)$ of norm $1$ cannot be very large on the data. As a counterexample, consider an inner product of the form $(\phi,\psi)=\int\phi(y)\psi(y)\ w(y)dy$, where the weighting function $w(y)$, though everywhere positive, is very small in at least one area where the true distribution $\rho(y)$ is not. Since the requirement that $\|g\|=1$ will effectively only constrain $g(y)$ in areas where $w(y)$ is comparatively large, the algorithm’s variables $b$ can make $L$ large by choosing test functions $g(y)$ not based on their correlation with $f(z)$ but on their amplitude in areas where $\rho$ is large but $w$ is small. Then the weight $w(y)$ should be such that the Radon–Nikodym derivative $\frac{d\rho}{dw}$ is bounded. We adopt an inner product that is fixed through stages of the procedure:
$\left<\phi,\psi\right>=\sum_{i}\phi\left(y^{0}_{i}\right)\psi\left(y^{0}_{i}\right),$ (7)
where $y^{0}_{i}$ is initially set to $x_{i}$ and updated between stages, to reflect the evolving distribution of the $\{y_{i}\}\sim\rho(y)$.
@@
In order to replace $G(y)$ by a smaller dimensional, orthogonal operator $Q_{y}(y)$ that spans the same effective range, we perform the reduced singular-value decomposition $G_{i}^{j}\stackrel{{\scriptstyle\text{def}}}{{=}}G^{j}\left(y^{0}_{i}\right)\approx\sum_{k=1}^{n_{y}}\sigma_{k}\ u^{k}_{i}\ v^{k}_{j}$, with $n_{y}\leq m_{y}$ chosen that the sum $\sum_{k=1}^{n_{y}}\left(\sigma_{k}\right)^{2}$ is larger that a fraction of the squared norm of $\left\|G\right\|^{2}=\sum_{i,j}\left(G_{i}^{j}\right)^{2}=\sum_{k=1}^{m_{y}}\left(\sigma_{k}\right)^{2}$, and we adopt $Q_{y}(y)=G(y)B^{y},\quad B^{y}_{jk}=\frac{1}{\sigma_{k}}v^{k}_{j}$ (Notice that, in particular, $Q_{y}^{k}\left(y^{0}_{i}\right)=u^{k}_{i}$.) With a fixed inner product, the matrix $B^{y}$ is computed only once per stage, when the $y^{0}$ are updated. The sense in which this $Q^{y}$ is orthogonal is not the conventional one: its columns represent the evaluation at arbitrary positions $\{y_{i}\}$ of a set of functions that are orthonormal under an inner product based on the $\{y^{0}_{i}\}$.
@@ 4 A flow-based methodology
The maximization over $a$ and $b$ can be carried out explicitly: they must align with the left and right first principal components of the $n_{z}\times n_{y}$ matrix $A^{hl}\stackrel{{\scriptstyle\text{def}}}{{=}}\sum_{i}Q^{h}_{z}\left(z_{i}\right)Q^{l}_{y}\left(y_{i}\right)$, and the penalty term is given by the first singular value $\sigma_{1}(y)$ of $A$. It follows that we can write the problem as a minimization over $y$ alone:
@@
$y_{i}^{n+1}=y_{i}^{n}-\eta^{n}\left[\frac{1}{n}\nabla_{y}\left.c\left(x_{i},y\right)\right|_{y_{i}^{n}}+\lambda\ a^{\prime}\left.\nabla_{y}A\right|_{y_{i}^{n}}b\right],\quad\left.\nabla_{y}A^{hl}\right|_{y_{i}^{n}}=Q^{h}_{z}\left(z_{i}\right)\sum_{j}\left.\nabla G^{j}(y)\right|_{y_{i}}B^{y}_{jl},$ (9)
for which all $\{y_{i}\}$ decouple, and $a$ and $b$ are updated in an alternate step.
@@
where $a$ and $b$ are the left and right $k_{\text{th}}$ principal components of $A$.
@@
The penalty term $\sigma_{1}(y)$ is not smooth at its arg-min $y=y^{*}$: for $y$ and $z$ to be independent, $A$ must vanish, and the first singular value $\sigma_{1}(y)$ of a matrix that depends smoothly on $y$ typically has corners where $A(y)$ vanishes (The simplest example is the $1\times 1$ matrix $A=y\in R$, whose only singular value $\sigma=|y|$ has a corner at $y=0$.) To address this, we square the penalty term: $\min_{\{y_{i}\}}L=\sum_{i=1}^{n}c\left(x_{i},y_{i}\right)+\lambda\ \sigma_{1}^{2}(y)$. There still remains one issue to address to make the methodology fully functional. Because every step of the algorithm brings down the largest singular value of $A(y)$, the first few of those singular values will tend to coalesce at convergence at a common value $\sigma_{*}\ll 1$. Then the derivatives of the penalty term with respect to the $\{y_{i}\}$ are not well defined, as they depend on the arbitrary choice of one pair among the various singular components $(a,b)$ associated to the singular value $\sigma_{*}$. In terms of test functions, more that one pair of functions $(f,g)$ have reached the threshold correlation $\sigma_{*}$. To address this, we modify the algorithm so that it tracks the first $K$ pairs $(a_{k},b_{k})$ of principal component of $A$, where $K=\min(\hbox{rank}(A),K_{max})$, with $K_{max}$ fixed by the user. Then we descend over $y$
@@
Notice that this extension carries little computational cost, since the $A(y)$ to differentiate is common to all the $\{\sigma_{k}\}$. If performed using reproducing Kernel Hilbert spaces, this extension could be thought as interpolating between COCO [27] and HSIC [34], where it is shown both that the penalty term vanishes if and only if $y$ and $z$ are independent, and that its empirical estimate converges to the population value at a rate $\frac{1}{\sqrt{n}}$.
@@ 5 Map inversion
Since we do not know $T(x,z)$ in closed form, it could appear that we can only invert it by learning $X(y,z)$ from its $n$ available samples $\{x_{i},y_{i},z_{i}\}$, for instance through kernel regression, nearest neighbor or neural networks. Yet we can do much better than that and obtain a closed form for $T^{-1}$, exploiting the fact that the penalization parameter $\lambda$ is large but finite. Since at convergence the gradient $\nabla_{y}L$ is zero, or at least sufficiently small to satisfy a termination criterion, we have $\nabla_{y_{i}}c\left(x_{i},y_{i}\right)+2\lambda\ \sum_{k=1}^{K}\sigma_{k}\nabla_{y}\sigma_{k}\Big|_{y_{i},z_{i}}=0$, which for the canonical quadratic cost in (2) yields $x_{i}=y_{i}+2\lambda\ \sum_{k=1}^{K}\sigma_{k}\nabla_{y}\sigma_{k}\Big|_{y_{i},z_{i}}$. In order to invert the map for arbitrary values of $y$ and $z$, we extend the validity of this expression and write
@@
The inversion formula in (11) provides us with a valuable bonus: it shows that the dependence of $x$ on $z$ that our algorithm has uncovered is mediated by the $K$ functions $\{f_{k}(z)\}$, so we have inadvertently performed factor extraction. These factors bring in insights about the mechanisms of the dependence of $x$ on $z$, while their gradient inform us of the sensitivity of $x$ with respect to changes in $z$. When we consider density estimation in Section 8, it will be useful to notice that we have access not only to $X(y,z)$ but also to its derivatives,